# Entanglement dynamics in intensity-dependent double Jaynes-Cummings model for squeezed coherent thermal states

### v6

@@ 7 Effects of Kerr-nonlinearity on the entanglement dynamics
The entanglement dynamics after adding Kerr-nonlinearity for IDDJCM are presented in Fig. 12. The addition of nonlinearity $\chi=0.1$ totally destroys the periodic nature of the entanglement for all the subsystem (blue curve in the plots). ESDs from the dynamics are removed for atom A-field a, atom A-atom b and field a-field b subsystems; however, in the case of atom A-atom B, ESDs become longer and peak heights get reduced significantly. As $\chi$ is increased to $0.3$, ESDs get longer in the dynamics of $C(t)$ and also peak which were present till then disappear. For other subsystems, amplitude of $N(t)$ for atom A-field a and filed-field entanglement decrease (green curve); however, for atom A-field b subsystem, $N(t)$ increases. For $\chi=0.7$, which is represented by the red curve in the plots, we observe that $C(t)$ begins to revive again, making the lengths of ESDs shorter. In atom A-field a and field a-field b subsystems, amplitude $N(t)$ decreases significantly, though there is no ESD present in the system. Like atom-atom entanglement, $N(t)$ increases for atom A-field b subsystem and length of ESDs decreases. Though, $N(t)$ decreases at the beginning of the dynamics, it amplitude increases with time. So, it can be concluded that there is a critical value $\chi$ for which $C(t)$ becomes minimum and beyond that value it increases again. As we increase the nonlinearity further to $\chi=1.0$, same pattern (black curve) is observed as we see for $\chi=0.7$ i.e., the atom-atom and atom A-field b entanglement increase while field-field and atom A-field a entanglement decrease further.
@@
The effects of $\chi$ on entanglement dynamics of different subsystems for SCTS in DJCM are presented in Fig. 13. For $\chi=0.1$, ESD in $C(t)$ increases considerably since the smaller peaks disappear from the dynamics (Fig. 13(a), blue curve). For the other subsystems, the dynamics of $N(t)$ does not change considerably for small values of $\chi$. As $\chi$ is increased to $0.3$, the amplitudes of $C(t)$ and $N(t)$ increase; however, ESDs are not removed from the dynamics. For other entanglements, $N(t)$ for atom A-field a and field a-field b decrease prominently with the increasing nonlinearity in the system. However, $N(t)$ for the atom A-field b subsystem increases and ESDs are removed from the dynamics. On further increasing the value of $\chi$ to $0.7$ and $1.0$, $C(t)$ increases very significantly with greater amplitude and ESDs are removed from the dynamics. Similar behaviour is observed for atom A-field b entanglement dynamics. On the other hand, for atom A-field a and field-field entanglement, $N(t)$ decreases prominently with these increased values of $\chi$. This shows that with the addition of $\chi$ to the system, entanglement is transferred in the atom-atom and atom A-field b subsystems from the atom A-field b and field a-field b subsystems.
@@ 8 Effects of detuning on the entanglement
Now, we discuss the effects of detuning on the entanglement dynamics of various subsystems. Detuning is defined as $\Delta=\omega-\nu$. In various studies, the effects of detuning on the dynamics of atom-field interaction have been investigated. The Hamiltonian of the atom-field system after adding detuning becomes
@@
$\hat{H}_{\text{det}}=\Delta\hat{\sigma}_{-}^{\text{A}}\hat{\sigma}_{+}^{\text{A}}+\lambda\left(\sqrt{\hat{N}_{\text{a}}}\,\hat{a}^{\dagger}\sigma_{-}^{\text{A}}+\hat{a}\sqrt{\hat{N}_{\text{a}}}\,\hat{\sigma}_{+}^{\text{A}}\right)+\Delta\hat{\sigma}_{-}^{\text{B}}\hat{\sigma}_{+}^{\text{B}}+\lambda\left(\sqrt{\hat{N}_{\text{b}}}\,\hat{b}^{\dagger}\hat{\sigma}_{-}^{\text{B}}+\hat{b}\sqrt{\hat{N}_{\text{b}}}\,\hat{\sigma}_{+}^{\text{B}}\right).$ (52)
$\hat{H^{\prime}}_{\text{det}}=\Delta\hat{\sigma}_{-}^{\text{A}}\hat{\sigma}_{+}^{\text{A}}+\lambda(\hat{a}^{\dagger}\hat{\sigma}_{-}^{\text{A}}+\hat{a}\hat{\sigma}_{+}^{\text{A}})+\Delta\hat{\sigma}_{-}^{\text{B}}\hat{\sigma}_{+}^{\text{B}}+\lambda(\hat{b}^{\dagger}\hat{\sigma}_{-}^{\text{B}}+\hat{b}\hat{\sigma}_{+}^{\text{B}}),$ (53)
To study the effects of detuning $\bar{n}_{c}=2,\bar{n}_{s}=1.0$ and $\bar{n}_{th}=0.1$ are chosen while the values of $\Delta$ are taken to be $2,5,10$. The effects of detuning on entanglements for SCTS in IDDJCM is plotted in Fig. 14. From the plots it is very much noticeable that $\Delta$ plays a significant role in removing ESDs from the system. For $\Delta=2$, length of ESDs get decreased in the dynamics of $C(t)$ (blue curve in Fig. 14(a)), by introducing new additional smaller peaks in the dynamics. Another important observation is that detuning decreases the peak heights of the entanglement pulses significantly with increasing time; however, the periodic nature of $C(t)$ is present to some extent. For atom A-field a entanglement, we see that the addition of $\Delta$ in the system totally destroys the periodic nature of the dynamics of $N(t)$. The amplitude of $N(t)$ is increased and all the ESDs are removed from the dynamics. The periodic-pulse like nature of $N(t)$ for atom A-field b is also gets destroyed by the addition of $\Delta$ to the system. For this smaller value of $\Delta$, the length of ESDs get smaller; however, they are not removed from the dynamics. In the case of field a-field b subsystem, it is observed that the addition of $\Delta=2$ removes all the ESDs from the dynamics of $N(t)$. Although, the periodic nature of $N(t)$ is present in the dynamics, the form of the periodic nature is distorted significantly.
For $\Delta=5$, the peak-heights of $C(t)$ come down further with smaller peaks appearing in some places of the dynamics. This makes the length of ESDs longer at some places and smaller at other places. Also, the entanglement pulses are destroyed with increasing $\Delta$. $N(t)$ for atom A-field a does not change considerably, while for atom A-field b, the amplitude of $N(t)$ increases and some of the ESDs get removed from the dynamics. For field a-field b subsystem, it is observed that the amplitude of $N(t)$ becomes almost half of the initial peak and the periodicity of the dynamics is totally lost.
As $\Delta$ is increased to $\Delta=10$, the length of initial ESD increases as a result of disappearance of other peaks from the dynamics. The dynamics becomes oscillatory for a short while, and again ESDs appear. So, we can see that detuning is not much effective to remove the ESDs from the dynamics of $C(t)$ and increases the value of atom-atom entanglement which is observed for SCS and G-L states.
The effects of $\Delta$ for SCTS in DJCM is depicted in Fig. 15. For $\Delta=2$, initial ESD is removed from the dynamics of $C(t)$; however, the peak following the second long ESD is lost (see black curve in Fig. 3(a) and blue curve in Fig. 15(a)), which makes the second ESD even longer. For all the other subsystems, the addition of $\Delta$ decreases the amplitude of $N(t)$. For the higher values of $\Delta=5,10$, though the amplitude of $C(t)$ increases more initially, the length of subsequent ESDs increases significantly. In the case of other subsystems, $N(t)$ is decreased further which makes ESD appear initially in field-field entanglement. From these plots it is evident that presence of thermal photons forbids entanglement transfer to atom-atom subsystem from other subsystems.
@@ 9 Conclusion
In this work, we have investigated the entanglement dynamics of the atom-atom, atom-field and field-field subsystems with squeezed coherent thermal states in the intensity-dependent double Jaynes-Cummings model (IDDJCM) and the standard double Jaynes-Cummings model (DJCM). It is noticed that the atom-atom and field-field entanglements remain almost unchanged with increasing number of squeezed photons while keeping coherent and thermal photons constant for IDDJCM. An array of entanglement pulses is formed with almost same ESDs and same peak amplitude. However, the atom-field entanglements are affected noticeably. For DJCM, all the dynamics change with increasing squeezed photons and length of ESDs also increases for atom-atom and atom A-field b subsystems.
The addition of photon exchange interaction between the cavities removes ESDs from field-field entanglement dynamics and from the atom-atom for large values of $\kappa$, for both IDDJCM and DJCM. The atom-field entanglements vanish for large values of $\kappa$. In the case of DJCM, formation of wave packets is observed for field-field subsystem.
Like photon exchange interaction, dipole-dipole interaction removes ESDs from the dynamics of $C(t)$ and $N(t)$ of field-field subsystem for high values of $g_{d}$ for both IDDJCM and DJCM. However, in this case, formation of wave packets is not observed. Ising interaction is effective to remove ESDs from field-field entanglement for IDDJCM; however, it is unable to remove ESDs from atom-atom and atom-field entanglements. For DJCM, the dynamics of $C(t)$ and $N(t)$ are not affected considerably by the addition of Ising interaction in the system.
Kerr-nonlinearity effectively removes ESDs from the all the entanglement dynamics. For IDDJCM, $C(t)$ and $N(t)$ for all the subsystems increase for higher values of $\chi$. In the case of DJCM also, $C(t)$ and $N(t)$ for atom-field subsystems increase for higher values of $\chi$; however, field-field entanglement decreases after a critical value of $\chi$. Detuning tries to remove ESDs from all the subsystems. ESDs vanish from field-field entanglement; however its peak-amplitude comes down with increasing value of $\Delta$. Also, for higher values of detuning, the $C(t)$ increases more and other entanglements decrease, which is an implications of entanglement transfer from other subsystems to atom-atom subsystem.

### v7

@@ 7 Effects of Kerr-nonlinearity on the entanglement dynamics
The entanglement dynamics after adding Kerr-nonlinearity for IDDJCM are presented in Fig. 11. The addition of nonlinearity $\chi=0.1$ totally destroys the periodic nature of the entanglement for all the subsystem (blue curve in the plots). ESDs from the dynamics are removed for atom A-field a, atom A-atom b and field a-field b subsystems; however, in the case of atom A-atom B, ESDs become longer and peak heights get reduced significantly. Addition of Kerr-nonlinearity in the system modifies the energy levels of the field depending on photon number, introducing non-uniform spacing between adjacent levels. As $\chi$ is increased to $0.3$, ESDs get longer in the dynamics of $C(t)$ and also peak which were present till then disappear. For other subsystems, amplitude of $N(t)$ for atom A-field a (green curve) and field-field (black curve) entanglement decrease; however, for atom A-field b subsystem, $N(t)$ increases. This happens because, the field b retain is retaining more structure or coherence under Kerr distortion, allowing it to maintain or enhance entanglement with atom A. For $\chi=0.7$, which is represented by the red curve in the plots, we observe that $C(t)$ begins to revive again, making the lengths of ESDs shorter. In atom A-field a and field a-field b subsystems, amplitude $N(t)$ decreases significantly, though there is no ESD present in the system. Like atom-atom entanglement, $N(t)$ increases for atom A-field b subsystem and length of ESDs decreases. Though, $N(t)$ decreases at the beginning of the dynamics, it amplitude increases with time. So, it can be concluded that there is a critical value $\chi$ for which $C(t)$ becomes minimum and beyond that value it increases again (see Appendix E for detail). As we increase the nonlinearity further to $\chi=1.0$, same pattern (black curve) is observed as we see for $\chi=0.7$ i.e., the atom-atom and atom A-field b entanglement increase while field-field and atom A-field a entanglement decrease further. The graphs indicate that the presence of Kerr nonlinearity favors the preservation of atom–atom and atom A–field b entanglement. This behavior suggests that Kerr nonlinearity redistributes quantum correlations among the subsystems, resulting in a competition for entanglement. Such redistribution is consistent with the principle of monogamy of entanglement, where the enhancement of entanglement in one pair limits its presence in others.
@@
The effects of Kerr-nonlinearity $(\chi)$ on the entanglement dynamics of various subsystems in the DJCM with squeezed coherent thermal states (SCTS) are depicted in Fig. 12. For $\chi=0.1$, the concurrence $C(t)$ for the atom–atom subsystem exhibits a significant increase in the duration of entanglement sudden death, as the smaller revival peaks disappear (Fig. 12(a), blue curve). In contrast, the entanglement dynamics of the other subsystems, as measured by the negativity $N(t)$, remain largely unaffected at this low Kerr strength. As $\chi$ increases to 0.3, the amplitudes of both $C(t)$ and $N(t)$ rise slightly; however, ESDs persist in the dynamics. Notably, atom A–field a and field a–field b entanglements decrease in amplitude, while atom A–field b entanglement increases, with ESDs being completely removed from its dynamics. This suggests the onset of entanglement redistribution driven by Kerr-induced spectral distortion.
Upon further increasing $\chi$ to 0.7 and 1.0, the concurrence $C(t)$ exhibits a substantial enhancement in amplitude, with complete removal of ESDs. A similar trend is observed for the atom A–field b entanglement, indicating a strong transfer of quantum correlations into these subsystems. Conversely, entanglement in the atom A–field a and field–field subsystems is significantly diminished.
These results indicate that in this case also, the Kerr-nonlinearity acts as a mechanism for entanglement redistribution within the system, favoring the atom–atom and atom A–field b subsystems at the expense of others. Moreover, the enhancement of $C(t)$ in the DJCM is observed to be more pronounced compared to the IDDJCM case.
@@ 8 Effects of detuning on the entanglement
Now, we discuss the effects of detuning on the entanglement dynamics of various subsystems. Detuning is defined as $\Delta=\nu-\omega$ which is difference between field frequency and atomic transition frequency. In various studies, the effects of detuning on the dynamics of atom-field interaction have been investigated. The Hamiltonian of the atom-field system in the rotating frame with $\Delta$ becomes[82]
@@
$\hat{H}_{\text{det}}=\Delta\hat{\sigma}_{-}^{\text{A}}\hat{\sigma}_{+}^{\text{A}}+\lambda\left(\sqrt{\hat{N}_{\text{a}}}\,\hat{a}^{\dagger}\hat{\sigma}_{-}^{\text{A}}+\hat{a}\sqrt{\hat{N}_{\text{a}}}\,\hat{\sigma}_{+}^{\text{A}}\right)+\Delta\hat{\sigma}_{-}^{\text{B}}\hat{\sigma}_{+}^{\text{B}}+\lambda\left(\sqrt{\hat{N}_{\text{b}}}\,\hat{b}^{\dagger}\hat{\sigma}_{-}^{\text{B}}+\hat{b}\sqrt{\hat{N}_{\text{b}}}\,\hat{\sigma}_{+}^{\text{B}}\right),$ (37)
$\hat{H^{\prime}}_{\text{det}}=\Delta\hat{\sigma}_{-}^{\text{A}}\hat{\sigma}_{+}^{\text{A}}+\lambda(\hat{a}^{\dagger}\hat{\sigma}_{-}^{\text{A}}+\hat{a}\hat{\sigma}_{+}^{\text{A}})+\Delta\hat{\sigma}_{-}^{\text{B}}\hat{\sigma}_{+}^{\text{B}}+\lambda(\hat{b}^{\dagger}\hat{\sigma}_{-}^{\text{B}}+\hat{b}\hat{\sigma}_{+}^{\text{B}}).$ (38)
To investigate the influence of detuning on the entanglement dynamics, we consider the average photon numbers $\bar{n}_{c}=2$, $\bar{n}_{s}=1.0$, and $\bar{n}_{th}=0.1$, while varying the detuning parameter as $\Delta=2,5,10$. The corresponding entanglement dynamics for the squeezed coherent thermal state (SCTS) in the intensity-dependent double Jaynes–Cummings model (IDDJCM) are illustrated in Fig. 13.
From the plots, it is evident that detuning $\Delta$ plays a crucial role in shaping the entanglement behavior and suppressing entanglement sudden deaths (ESDs). For a moderate detuning value $\Delta=2$, the atom–atom entanglement $C(t)$ (Fig. 13(a), blue curve) shows a significant reduction in the duration of ESDs due to the emergence of additional smaller revival peaks. However, while these smaller revivals reduce ESD duration, the overall peak heights diminish with time, and the dynamics exhibit partial loss of periodicity.
In the case of atom A–field a entanglement $N(t)$, detuning destroys the periodic pulse structure entirely and enhances the amplitude, effectively eliminating ESDs. A similar behavior is observed for the atom A–field b entanglement, where $\Delta$ breaks the periodicity and shortens the ESD intervals, though not entirely removing them. For the field a–field b subsystem, $\Delta=2$ successfully removes all ESDs, although the resulting dynamics become less regular and the periodic structure becomes distorted.
As detuning increases to $\Delta=5$, the peak height of $C(t)$ further diminishes, while irregular smaller peaks appear sporadically, leading to inconsistent ESD durations across the time evolution. In this case, atom A–field a entanglement remains relatively unaffected, whereas atom A–field b entanglement experiences increased amplitude and partial removal of ESDs. For field–field entanglement, $N(t)$ loses its periodic nature entirely and its amplitude is significantly reduced.
For large detuning, $\Delta=10$, the system displays more prominent oscillatory suppression. The atom–atom entanglement initially shows longer ESDs due to the disappearance of revival peaks. After a short revival, ESDs reappear, indicating that large detuning does not effectively prevent ESD in the long run. These findings are consistent with earlier results obtained for squeezed coherent and Glauber–Lachs states [59], where detuning enhances atom–atom entanglement but limits energy and information exchange across subsystems.
The effect of detuning on SCTS within the standard DJCM framework is shown in Fig. 14. For $\Delta=2$, the initial ESD in $C(t)$ is removed, yet the subsequent peak is suppressed, causing the next ESD to lengthen significantly (compare blue curves in Fig. 2(a) and Fig. 14(a)). For all atom–field and field–field subsystems, $N(t)$ decreases notably, indicating reduced entanglement sharing across the subsystems.
At higher detuning values $\Delta=5$ and $10$, the amplitude of $C(t)$ initially increases but the corresponding ESDs become much longer, highlighting an accumulation of entanglement in the atom–atom subsystem at the expense of the others. Meanwhile, the atom–field and field–field entanglements decay further, with ESD appearing early in the field–field entanglement dynamics. This behavior suggests that increasing detuning transfers entanglement to certain subsystems (particularly atom–atom), from other subsystems. This is consistent with the previous results in [59]. As a result, detuning serves as a control knob for selectively enhancing or suppressing subsystem entanglements, which is valuable in designing robust quantum information protocols.
@@ 9 Conclusion
In this work, we have systematically investigated the entanglement dynamics of atom–atom, atom–field, and field–field subsystems in the presence of squeezed coherent thermal states (SCTS) within both the intensity-dependent double Jaynes–Cummings model (IDDJCM) and the standard double Jaynes–Cummings model (DJCM). Our analysis reveals distinct behaviors in entanglement evolution depending on the model type and the nature of interactions introduced. In the IDDJCM, atom–atom and field–field entanglements remain largely unaffected by variations in the average number of squeezed photons when coherent and thermal photon numbers are held constant, whereas atom–field entanglements show noticeable sensitivity. In contrast, DJCM dynamics respond more strongly to increased squeezing, with entanglement sudden death (ESD) becoming more pronounced in the atom–atom and atom A–field b subsystems.
The inclusion of photon exchange interaction between the cavities plays a crucial role in removing ESDs from the dynamics, particularly in the atom–atom and field–field subsystems, for sufficiently large values of the interaction parameter $\kappa$, in both DJCM and IDDJCM. However, this comes at the cost of diminishing atom–field entanglements. Notably, in DJCM, the field–field entanglement exhibits wave-packet-like structures under strong photon hopping, which are absent in IDDJCM. Similarly, the introduction of dipole–dipole interaction ($g_{d}$) effectively eliminates ESDs from the atom–atom and field–field subsystems in both models, although no wave packet formation is observed in this case. On the other hand, Ising-type interaction exhibits limited utility: while it can remove ESDs from field–field entanglement in IDDJCM, it fails to do so for atom–atom and atom–field subsystems, and has minimal effect on DJCM dynamics.
Kerr nonlinearity further enriches the entanglement landscape. In both models, it enhances $C(t)$ and $N(t)$ for several subsystems and efficiently removes ESDs, particularly in IDDJCM. However, for DJCM, field–field entanglement begins to decline beyond a critical Kerr strength $\chi$, indicating a nonlinear redistribution of entanglement. Detuning also contributes to controlling entanglement dynamics by suppressing ESDs and promoting robustness. Increasing the detuning parameter $\Delta$ leads to a rise in atom–atom entanglement while attenuating the strength of atom–field and field–field correlations, suggesting a transfer of entanglement into the atom–atom channel.
From an application perspective, the study of SCTS in the context of the IDDJCM offers valuable insights for quantum information science and quantum technology. The photon-number-dependent coupling in IDDJCM models nonlinear atom–field interactions relevant to real-world platforms such as superconducting qubits, trapped ions, and cavity QED systems. The mixed-state character of SCTS captures realistic environmental effects by combining coherence, squeezing, and thermal noise. These results hold particular relevance for the development of robust quantum communication protocols, fault-tolerant quantum gates, and long-lived quantum memories. By understanding how entanglement behaves under various interactions and noisy conditions, this study contributes to the ongoing effort to design quantum systems with tunable, stable, and controllable entanglement—essential ingredients for scalable quantum technologies.